OpenStudy (anonymous):

Arithmetic Sequence: Given the first element 19/7 and the common difference -2/7 and the Sn(arithmetic series) 13. Find the value of the nth element and the value of N

OpenStudy (anonymous):

\[a _{1}=19/7\]

OpenStudy (anonymous):

\[\huge s _{n}=13\]

OpenStudy (anonymous):

\[\huge a _{n}\] is the nth element.

OpenStudy (anonymous):

what do u think?

OpenStudy (anonymous):

formulas: \[\huge s _{n}=(N/2)(a _{1}+a _{n})\]

OpenStudy (anonymous):

\[\huge a _{n}= a _{1} + (N-1)d\]

OpenStudy (anonymous):

u had to find nth element so will apply this formula \[_{Sn}=n/2(2a+(n-1)d)\]

OpenStudy (anonymous):

ok. then what? :(

OpenStudy (anonymous):

this is part of our quiz and i didnt get the correct answer. thats why im asking for help because our exam is tomorrow

OpenStudy (anonymous):

then after u will get a quadratic equation in terms of n. So u had solve for n.

OpenStudy (anonymous):

\[(−34N2+490N−3094)/119\]

OpenStudy (anonymous):

\[n^{2}-20n+91\] will get this equation

OpenStudy (anonymous):

how did you get that? :((

OpenStudy (anonymous):

please show me the soln. :(

OpenStudy (anonymous):

oh. shoot.i wrote something wrong in the given. a sub 1 should be equal to 19/17

OpenStudy (anonymous):